OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. (2x+12)^2/3-3=13

hero (hero):

(2x+12)^2/3 = 16 (2x+12)^2 = 48 2x+12 = pm sqrt{48} 2x+12 = pm 4 sqrt{3} 2x = pm 4sqrt{3} - 12 x = ((pm 4sqrt{3})/2) - 6 I don't believe there are extraneous solutions here Extraneous solutions would be values of x that if you put it back into the original equation, you don't get 16